# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v1

@@ Appendix A Social Force Model
Sliding friction coefficient ($\kappa$) $2.4\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
Interaction strength ($A_{i}$) $2\times 10^{3}\,\text{N}$
Interaction range ($B_{i}$) $0.08\,\text{m}$
Interaction strength ($C_{i}$) $2\times 10^{3}\,\text{N}$
Interaction range ($D_{i}$) $0.08\,\text{m}$
Mass $80\,\text{kg}$
Relaxation time ($\tau$) $0.5\,\text{s}$
Pedestrian radius $0.2\,\text{m}$
Initial velocity in $x$-direction ($v_{x}$) $0\,\text{m/s}$
Initial velocity in $y$-direction ($v_{y}$) $0\,\text{m/s}$
Obstacle width $3.6\,\text{m}$
Obstacle height $3.6\,\text{m}$
$x$-coordinate of center $25.8\,\text{m}$
$y$-coordinate of center $1.8\,\text{m}$
@@ Appendix B Initial Conditions of Microscopic Distributions
Here, we describe the initial conditions that we considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested. We considered zero initial velocities for all pedestrians, while we used the following distributions for initializing the position $\bm{x}_{i}=(x_{i},y_{i})\in\Omega$ of the $i$-th pedestrian.
  • Positions are sampled from a two dimensional normal distribution with mean $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
@@
  • Uniform density distribution: Positions are sampled uniformly over the domain $\Omega$
    $p(x_{i},y_{i})=\mathcal{U}\left([x_{min},x_{max}]\times[y_{min},y_{max}]\right).$ (B.2)
  • Cosine distribution;
    $p(x_{i},y_{i})=\frac{1}{\pi\sigma_{x}\sigma_{y}}\cos\left(\frac{x_{i}-\mu_{x}}{\sigma_{x}}\right)\cos\left(\frac{y_{i}-\mu_{y}}{\sigma_{y}}\right).$ (B.3)
  • Piecewise linear density distribution:
    To initialize particle positions in a rectangular domain $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$, we assume a uniform distribution in $x$ and a non-uniform, piecewise linear distribution in $y$. This approach allows control over the particle concentration along the vertical axis.
    Particles are sampled uniformly in the $x$-direction over the interval $x_{\min},x_{\max}]$. In the $y$-direction, particle positions are sampled from a custom piecewise linear distribution designed to increase density near the top and bottom boundaries. Specifically, we use a triangular distribution defined over the interval $[0,12]$ by
    $p(y_{i})=\begin{cases}\frac{2}{12}\left(1-\frac{y_{i}}{6}\right),&0\leq y_{i}\leq 6,\\ \frac{2}{12}\left(\frac{y_{i}-6}{6}\right),&6<y_{i}\leq 12,\end{cases}$ (B.4)
    which has peaks at $y=0$ and $y=12$, and a minimum at $y=6$.
    To sample $y$-coordinates from this distribution, we use inverse transform sampling. First, a uniform random number $u\sim\mathcal{U}(0,1)$ is generated. If $u\leq 0.5$, the corresponding $y$ value is computed by solving $u=\frac{y}{12}(2-\frac{y}{6})$, yielding
    $y=6\left(1-\sqrt{1-2u}\right).$ (B.5)
    If $u>0.5$, we solve $u=\frac{1}{2}+\frac{(y-6)^{2}}{72}$, resulting in
    $y=6+\sqrt{72(u-0.5)}.$ (B.6)
    This method produces particle positions that are uniformly distributed along the horizontal axis and more concentrated near the top and bottom edges along the vertical axis.
For generating the training/testing data set, we used a wide range of initial conditions, based on the described distributions;in particular, for the gaussian sampling, the means and standard deviations were selected relative to the domain size to ensure spatial coverage while avoiding boundary artifacts. Specifically, the means were placed at the geometric center or fractional positions within the domain (e.g., $0.5$ or $0.25$ of the domain width), while standard deviations were set as fractions of the domain dimensions (e.g., one-fifth or one-half of the width or height). This ensures that the pedestrian positions lie within the domain, while allowing for spatial clustering or multimodal patterns. The parameters were chosen heuristically to reflect plausible crowd configurations and to introduce variety in the training and testing datasets. the ones used to construct the training are shown in Table B.1, while the ones for the testing data set are shown in Table B.2.
@@
Table B.1: Microscopic distribution configurations and parameters for training the LSTM and MVAR models.
Distribution Type Parameters
Data set 1: Gaussian $\begin{aligned} &\mu_{x}=7.5,\;\mu_{y}=5.0,\\ &\sigma_{x}=2.2,\;\sigma_{y}=2.0\end{aligned}$
Data set 2: Uniform $\mathcal{U}\left([2,15]\times[3,6]\right)$
Data set 3: Gaussian $\begin{aligned} &\mu_{x}=8.4,\;\mu_{y}=4.0,\\ &\sigma_{x}=1.8,\;\sigma_{y}=1.3\end{aligned}$
Data set 4: Gaussian $\begin{aligned} &\mu_{x}=12.5,\;\mu_{y}=4.0,\\ &\sigma_{x}=7,\;\sigma_{y}=1.1\end{aligned}$
Data set 5: Double Gaussian $\begin{aligned} &\mu_{x1}=17.3,\;\mu_{x2}=13.6,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.7,\\ &\sigma_{y}=1.2;\;\mu_{y}=7.0\end{aligned}$
Data set 6: Double Gaussian $\begin{aligned} &\mu_{x1}=19.7,\;\mu_{x2}=13.7,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.62,\\ &\sigma_{y}=4.0;\;\mu_{y}=2.33\end{aligned}$
Data set 7: Double Gaussian $\begin{aligned} &\mu_{x1}=32.2,\;\mu_{x2}=27.75,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.35,\\ &\sigma_{y}=4.0;\;\mu_{y}=1.0\end{aligned}$
Data set 8: Piecewise Linear $x_{i}\sim\mathcal{U}\left(30,47\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
Data set 9: Piecewise Linear $x_{i}\sim\mathcal{U}\left(15,35\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
Data set 10: Cosine $\begin{aligned} &\mu_{x}=25,\;\mu_{y}=3.5,\\ &L_{x}=10,\;L_{y}=2.5\end{aligned}$
@@
Table B.2: Microscopic distribution configurations and parameters for testing the LSTM and MVAR models.
Configuration Parameters
Dataset 1: Double Gaussian $\begin{aligned} &\mu_{x1}=41.2,\;\mu_{x2}=33.7,\\ &\sigma_{x1}=\sigma_{x2}=0.5,\\ &\sigma_{y}=2.0,\;\mu_{y}=9.0\end{aligned}$
Dataset 2: Double Gaussian $\begin{aligned} &\mu_{x1}=11.7,\;\mu_{x2}=5.2,\\ &\sigma_{x1}=\sigma_{x2}=0.4,\\ &\sigma_{y}=1.0,\;\mu_{y}=9.0\end{aligned}$
Dataset 3: Double Gaussian $\begin{aligned} &\mu_{x1}=42.5,\;\mu_{x2}=37.5,\\ &\sigma_{x1}=\sigma_{x2}=0.3,\\ &\sigma_{y}=2.0,\;\mu_{y}=5.0\end{aligned}$
Dataset 4: Gaussian $\begin{aligned} &\mu_{x}=36.0,\;\mu_{y}=4.0,\\ &\sigma_{x}=2.4,\;\sigma_{y}=1.5\end{aligned}$
Dataset 5: Gaussian $\begin{aligned} &\mu_{x}=23.0,\;\mu_{y}=8.5,\\ &\sigma_{x}=2.0,\;\sigma_{y}=2.5\end{aligned}$
Dataset 6:Gaussian $\begin{aligned} &\mu_{x}=7.85,\;\mu_{y}=8.0,\\ &\sigma_{x}=4.3,\;\sigma_{y}=2.0\end{aligned}$
Dataset 7: Cosine $\begin{aligned} &\mu_{x}=40,\;\mu_{y}=8.0,\\ &L_{x}=5,\;L_{y}=2.0\end{aligned}$
Dataset 8: Uniform $\mathcal{U}\left([10,20]\times[4,9]\right)$
Dataset 9: Piecewise Linear $x_{i}\sim\mathcal{U}\left(30,44\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
Dataset 10: Cosine $\begin{aligned} &\mu_{x}=39,\;\mu_{y}=6.5,\\ &L_{x}=7,\;L_{y}=2.5\end{aligned}$
Appendix C Kernel Density Estimation
@@
Here, we provide additional details for the first step of the proposed framework, regarding the extraction of continuous density profiles from discrete pedestrians’ positions. As already discussed in Section 2.2.1, the Kernel Density Estimation (KDE) is used for this task, following the general form in Eq. (6). Here, we elaborate on the kernel selection, bandwidth tuning, and adjustments for boundaries and obstacles.
@@
The choice of kernel function directly influences the smoothness and continuity of the density estimate. Common options include the uniform, triangular, Epanechnikov, and Gaussian kernels [95, 96]. The Epanechnikov kernel is theoretically optimal for minimizing mean integrated squared error (MISE) [97], but its bounded support can limit flexibility in practice. In contrast, the Gaussian kernel is widely preferred due to its smoothness, infinite support, and analytical tractability. Its multivariate form is given by:
@@
$h_{i}=\left(\frac{4}{N(d+2)}\right)^{1/(d+4)}\sigma_{i}\quad(i=x,y).$ (C.2)
where $N$ is the number of data points, $d$ is the number of spatial dimensions, in this case $d=2$ and $\sigma_{i}$ is the standard deviations of the pedestrian positions. However, in our case, Silverman’s rule resulted in wide bandwidths, which did not allow us to consider sufficient detail in the density profiles. We thus tuned $h_{x}/h_{y}$ to the values reported in Section 2.2.1.
@@
while the support in the $y$-direction remains unbounded, i.e., $\Omega_{y}=(-\infty,\infty)$. To enforce continuity and reduce boundary-induced artifacts in the $x$-direction, we further augment the observed positions with “ghost” particles near the domain edges. Let $\epsilon=\Delta x$ define a threshold for proximity to the boundary. For each position $\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at

### v2

@@ Appendix A Social Force Model
Corridor length ($L_{x}$) $48\,\text{m}$ Sliding friction coefficient ($\kappa$) $2.4\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
Corridor width ($L_{y}$) $12\,\text{m}$ Interaction strength ($A_{i}$) $2\times 10^{3}\,\text{N}$
Obstacle height $3.6\,\text{m}$ Interaction range ($B_{i}$) $0.08\,\text{m}$
Obstacle width $3.6\,\text{m}$ Interaction strength ($C_{i}$) $2\times 10^{3}\,\text{N}$
Obstacle $x$-coord. center $25.8\,\text{m}$ Interaction range ($D_{i}$) $0.08\,\text{m}$
Obstacle $y$-coord. center $1.8\,\text{m}$ Pedestrian mass ($m_{i}$) $80\,\text{kg}$
Initial velocity in $x$-direction ($v_{x}$) $0\,\text{m/s}$ Relaxation time ($\tau$) $0.5\,\text{s}$
Initial velocity in $y$-direction ($v_{y}$) $0\,\text{m/s}$ Pedestrian radius ($r_{i}$) $0.2\,\text{m}$
Simulation time $250\,\text{s}$ Time Step ($\delta_{t}$) $2.5\times 10^{-2}\,\text{s}$
For our simulations, the model parameters were calibrated based on established values from previous studies [1, 70]. These works describe pedestrians with a standard average mass of 80kg80\ kg, for which a relaxation time $\tau=0.5\,\text{s}$ provides adequate behavior under normal conditions. The complete set of parameters governing repulsive potentials, including pedestrian-pedestrian and pedestrian-obstacle interactions, along with the simulation environment specifications, are detailed in Table A.1. This parameter selection, which is uniform for all pedestrians, ensures consistency with validated bottleneck evacuation scenarios. Although individual variations exist in reality, using a common set of parameters is a standard approach in the literature [70] to maintain a manageable calibration process while enabling the reliable simulation of diverse crowd dynamics, including the present case study.
@@ Appendix B Initial Conditions of Microscopic Distributions
Here, we describe the initial conditions considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested. We considered zero initial velocities for all pedestrians, while we used the following distributions for initializing the position $\bm{x}_{i}=(x_{i},y_{i})\in\Omega$ of the $i$-th pedestrian. Positions are sampled from:
  • Uniform distributions over the domain $x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\subset\Omega$:
    $p(x_{i},y_{i})=\mathcal{U}\left([x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\right).$ (B.1)
  • Two dimensional normal distributions with means $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
@@
  • “Double bell-shaped” independent joint distributions, which are comprised by normal distributions along the yy–axis with mean $\mu_{y}$ and standard deviation $\sigma_{y}$, and a mixture of normal distributions along the xx–axis with means $(\mu_{x_{1}},\mu_{x_{2}})$ and standard deviations $(\sigma_{x_{1}},\sigma_{x_{2}})$:
    $p(x_{i},y_{i})$ $=\frac{1}{\sqrt{2\pi}\,\sigma_{y}}\,\exp\!\left(-\frac{(y-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right)$ (B.3)
    $\quad\times\frac{1}{2}\left[\frac{1}{\sqrt{2\pi}\,\sigma_{x_{1}}}\exp\!\left(-\frac{(x-\mu_{x_{1}})^{2}}{2\sigma_{x_{1}}^{2}}\right)+\frac{1}{\sqrt{2\pi}\,\sigma_{x_{2}}}\exp\!\left(-\frac{(x-\mu_{x_{2}})^{2}}{2\sigma_{x_{2}}^{2}}\right)\right].$
  • Cosine distributions centered at $(c_{x},c_{y})$ with scale parameters $(s_{x},s_{y})$:
    $p(x_{i},y_{i})=\frac{1}{\pi s_{x}s_{y}}\cos\left(\frac{x_{i}-c_{x}}{s_{x}}\right)\cos\left(\frac{y_{i}-c_{y}}{s_{y}}\right).$ (B.4)
The training and testing data sets were generated from $C=20$ initial conditions, with 5 samples drawn from each of the four distributions described above. To ensure plausible and varied crowd configurations, the parameters for these distributions were selected heuristically to provide a broad spatial coverage while mitigating boundary artifacts. Furthermore, during the sampling process, pedestrian positions were resampled if they were within a critical distance equal to $2r_{i}$ of an existing pedestrian or a wall/obstacle to prevent initial overlaps and collisions. The specific parameter configurations used for the training and testing data sets are detailed Table B.1 and Table B.2, respectively.
@@
Table B.1: Microscopic initial distribution configurations and parameters for constructing the training dataset.
Case c Distribution Type Parameters
1 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=15,y_{\text{min}}=3,y_{\text{max}}=9$
2 Uniform, Eq. (B.1) $x_{\text{min}}=33,x_{\text{max}}=46,y_{\text{min}}=3,y_{\text{max}}=9$
3 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=20,y_{\text{min}}=4,y_{\text{max}}=8$
4 Uniform, Eq. (B.1) $x_{\text{min}}=28,x_{\text{max}}=46,y_{\text{max}}=4,y_{\text{max}}=8$
5 Uniform, Eq. (B.1) $x_{\text{min}}=3,x_{\text{max}}=17,y_{\text{max}}=4,y_{\text{max}}=8$
6 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
7 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
9 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=5,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=8,\;\mu_{y}=7,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=0.8$
12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=28,\;\mu_{y}=6,\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=32,\;\sigma_{x_{2}}=1.1$
14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=36,\;\mu_{y}=6,\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=40,\;\sigma_{x_{2}}=1.2$
15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=9,\;\mu_{y}=6,\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=1$
16 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
17 Cosine, Eq. (B.4) $c_{x}=20,\;c_{y}=8,s_{x}=10,\;s_{y}=3$
18 Cosine, Eq. (B.4) $c_{x}=30,\;c_{y}=7,s_{x}=10,\;s_{y}=3$
19 Cosine, Eq. (B.4) $c_{x}=40,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
20 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=9,\;s_{y}=3$
@@
Table B.2: Microscopic initial distribution configurations and parameters for constructing the testing dataset.
Case c Distribution Type Parameters
1 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=16,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
2 Uniform, Eq. (B.1) $x_{\text{min}}=32,\;x_{\text{max}}=46,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
3 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=19,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
4 Uniform, Eq. (B.1) $x_{\text{min}}=29,\;x_{\text{max}}=46,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
5 Uniform, Eq. (B.1) $x_{\text{min}}=4,\;x_{\text{max}}=17,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
6 Gaussian, Eq. (B.2) $\mu_{x}=12,\;\mu_{y}=7,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
7 Gaussian, Eq. (B.2) $\mu_{x}=11,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=2$
8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
9 Gaussian, Eq. (B.2) $\mu_{x}=39,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=7,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=7,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.1,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=0.9$
12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=29,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=33,\;\sigma_{x_{2}}=1.1$
14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=37,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=41,\;\sigma_{x_{2}}=1.2$
15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\;\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=1$
16 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=4,\;s_{x}=8,\;s_{y}=3$
17 Cosine, Eq. (B.4) $c_{x}=25,\;c_{y}=6,\;s_{x}=10,\;s_{y}=3$
18 Cosine, Eq. (B.4) $c_{x}=35,\;c_{y}=5,\;s_{x}=10,\;s_{y}=3$
19 Cosine, Eq. (B.4) $c_{x}=45,\;c_{y}=7,\;s_{x}=5,\;s_{y}=3$
20 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=6,\;s_{x}=7,\;s_{y}=3$
Appendix C Density fields extraction via Kernel Density Estimation
@@
Here, we provide additional details for the first step of the proposed framework, regarding the extraction of continuous density profiles from discrete pedestrians’ positions. As already discussed in Section 2.2.1, we used KDE for this task, following the general form in Eq. (6); here, we elaborate on the kernel selection, bandwidth tuning, and adjustments for boundaries and obstacles.
@@
The choice of kernel function directly influences the smoothness and continuity of the density estimate. Common options include the uniform, triangular, Epanechnikov, and Gaussian kernels [106, 107]. The Epanechnikov kernel is theoretically optimal for minimizing mean integrated squared error [108], but its bounded support can limit flexibility in practice. In contrast, the Gaussian kernel is widely preferred due to its smoothness, infinite support, and analytical tractability. Its multivariate form is given by:
@@
$h_{i}=\sigma_{i}\left(\frac{4}{N(d+2)}\right)^{1/(d+4)},$ (C.2)
where $N$ is the number of data points, $d$ is the number of spatial dimensions, in this case $d=2$, and $\sigma_{i}$ is the standard deviation of the pedestrian positions along $i=x,y$ spatial directions. However, in our case, Silverman’s rule resulted in wide bandwidths, which did not allow us to consider sufficient detail in the density profiles. We thus tuned $h_{x}/h_{y}$ to the values reported in Section 2.2.1.
@@
while the support in the $y$-direction remains unbounded. To enforce continuity and reduce boundary-induced artifacts in the $x$-direction, we further augment the observed positions with “ghost” particles near the domain edges. Let $\epsilon=\Delta x$ define a threshold for proximity to the boundary. For each position $\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at